If x is the sum of seven consecutive odd integers beginning with 3 and y is the sum of seven consecutive odd integers beginning with 5, then y-x equals
A 7 b 8 c 12 d 14

Answer by checkley77(12844)

x=3+5+7+9+11+13+15=63
y=5+7+9+11+13+15+17=77
77-63=14
Answer D) 14.